OpenStudy (anonymous):

(y^2+6y-17)^((1)/(3))-(y^2+3y-5)^((1)/(3))=0

hero (hero):

\[\frac{\sqrt[3]{y^2 + 6y - 17}}{\sqrt[3]{y^2 + 3y - 5}} = 0\] \[\sqrt[3]\frac{{y^2 + 6y - 17}}{{y^2 + 3y - 5}} = 0\] \[\left(\sqrt[3]\frac{{y^2 + 6y - 17}}{{y^2 + 3y - 5}}\right)^3 = 0\] \[\frac{{y^2 + 6y - 17}}{{y^2 + 3y - 5}}= 0\] \[y^2 + 6y - 17 = 0\] Finish solving by completing the square.

hero (hero):

@salce08